A 5000 kg open train car is rolling on frictionless rails at 22 m/s when it starts pouring rain. A few minutes later, the car's speed is 20 m/s. What mass of water has collected in the car?
Answer #1

Solution: Mass of car = M1 = 5000kg.

Initial speed of car = Vi = 22m/s.

final speed of car with water = Vf =20m/s.

Let mass of car with water =M2

  Since it is rolling on frictionless rails so, external force exerted on car is zero.

Hence momentum is conserve.

initial momentum = final momentum.

M1*Vi = M2*V2

5000*22 = M2*20

So, M2 = 5500kg

Hence mass of collected water = 5500 - 5000 = 500kg.